OpenStudy (anonymous):

The length of one side of a square is 86 cm. What is the area of the square?

OpenStudy (whpalmer4):

\[A = s^2\]where \(s\) is the side length of the square.

OpenStudy (anonymous):

that is not one of the answers

OpenStudy (whpalmer4):

That's the formula you need to use to find the area of the square.

OpenStudy (anonymous):

oh can you help me im not good with math really

OpenStudy (whpalmer4):

\[A = s^2\]\[A = s*s\]\[A = (86\text{ cm})*(86\text{ cm}) = \text{__________} \text{ cm}^2\]

OpenStudy (whpalmer4):

Multiply 86*86 and fill in the blank.

OpenStudy (anonymous):

7396

OpenStudy (whpalmer4):

Very good! So the answer is \(7396 \text{ cm}^2\) or \(7396 \text{ square centimeters}\)

OpenStudy (anonymous):

thanks i would give you a medal but i dont know how

OpenStudy (whpalmer4):

click the blue Best Response button